OpenStudy (anonymous):

what is the minimum value of sin(x) -1

OpenStudy (anonymous):

a) 1 b) -1 c) 0 d) -2

OpenStudy (anonymous):

please hurry..

hartnn (hartnn):

mnimum value of sin x is -1 so min value of sin x-1 = -1-1 =... ?

OpenStudy (anonymous):

its 0

hartnn (hartnn):

no, -1-1 is not 0

hartnn (hartnn):

=-(1+1) =.... ?

OpenStudy (anonymous):

ohokay so whats the asnwr

OpenStudy (hba):

What is -(1+1)

hartnn (hartnn):

you don't know whats 1+1 =... ?

OpenStudy (shubhamsrg):

hahahahahaha

OpenStudy (hba):

LoL

OpenStudy (aravindg):

lols

OpenStudy (unklerhaukus):